OpenStudy (anonymous):

could someone please help me i will reward a medal !!!!!!!! Sam wanted to leave an 18% tip on his $24 bill, what is the total amount that he has to pay?

OpenStudy (kc_kennylau):

What's the tip that he paid?

OpenStudy (anonymous):

0.18*24 = ? 24 + ? = total amount

OpenStudy (anonymous):

is the answer 456 ?

OpenStudy (kc_kennylau):

Nope

OpenStudy (kc_kennylau):

Mind showing us how you got 456?

OpenStudy (anonymous):

456 is quite a bit, don't you think? :p Especially if you were attending to pay a $24 bill. That must be one heck of a tip. 18/100 * 24 do that instead.

OpenStudy (anonymous):

so 18/100 would be 0.18 then x 24 = 4.32

OpenStudy (kc_kennylau):

Of course 18/100 is 0.18

OpenStudy (kc_kennylau):

Okay now add the tip to the original amount of money

OpenStudy (anonymous):

so 0.18+24=24.18

OpenStudy (kc_kennylau):

The tip is not 0.18

OpenStudy (anonymous):

?? idk if thats right or wrong

OpenStudy (kc_kennylau):

18% is the same as 0.18 which is the same as 18/100

OpenStudy (kc_kennylau):

But 18% of 24 is not 0.18

OpenStudy (anonymous):

oh sorru the tip if 4.32

OpenStudy (anonymous):

Yes, now add that tip onto the total amount of $24.

OpenStudy (anonymous):

k so 28.32

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

Good job! :P

OpenStudy (anonymous):

thank you for helping me :)))))))))))))))

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

really means alot to both of you @iambatman @kc_kennylau thanks so much!

OpenStudy (kc_kennylau):

You're welcome :)

OpenStudy (anonymous):

1.18 * 24=28.32

OpenStudy (anonymous):

$5.00

OpenStudy (kc_kennylau):

This is the most random answer I've ever seen

OpenStudy (kc_kennylau):

How on earth did you get this lol I'd love to see the steps